OpenStudy (angelax0):

A line passes through the point (8,−8) and has a slope of −3/4 . Write an equation in slope-intercept form for this line.

OpenStudy (rcp031):

x=8, y=-8, values of ordered pair (8,−8) We know that slope-intercept form is y=mx+c here x=8, y=-8, m=slope=rise/run =−3/4 & c=constant Find value of C -8=(-3/4)*8+c c=-6+8=-2 hence,equation is y=-3x/4-2 check put x=8 u l'll get y=-8

OpenStudy (angelax0):

do you mean y=-3/4x-2 ?

OpenStudy (rcp031):

Sorry typos problem Final equation will be y=(-3x/4)-2 here m=-3/4 & c=-2 sorry for typos.

OpenStudy (angelax0):

thank you

OpenStudy (rcp031):

NP:)